Multiplicity of Unstable Nontrivial Solutions in Semilinear Elliptic Indefinite Weight Boundary Value Problems

Bongsoo Ko Cheju National University

We discuss the existence and multiplicity of unstable nontrivial solutions about the following problems: −Δu = λg(x)f(u) in Ω, (1 − α)\frac{∂u}{∂n} + αu = 0 on ∂Ω. Here, Ω is a bounded open region in R^{n}, with the smooth boundary ∂Ω.